OpenStudy (anonymous):

solve by factoring x^2+8x=20? Can anyone show me how to do this, i'm struggling!

OpenStudy (anonymous):

i know x=4, but i don't remember how to work this?

OpenStudy (anonymous):

good to know you'restruggling eith a question...... x^2+8x-20=0 x^2+10x -2x-20=0 x(x+10)-2(x+10)=0 (x+10)(x-2)+0 x=-10 or 2

OpenStudy (anonymous):

x=2*

OpenStudy (anonymous):

its=0

OpenStudy (anonymous):

not+0

OpenStudy (anonymous):

Ohh okay, i see what i was doing wrong, thanks(: Would you mind helping with one more?

OpenStudy (anonymous):

@Meinme

OpenStudy (anonymous):

yeah feel free

OpenStudy (anonymous):

Okay, this one is: Solve by quadratic formula 5x^2 + 2x - 7 = 0

OpenStudy (anonymous):

b^2-4ac= 4+140=144 by quadratic formula we have 

OpenStudy (anonymous):

ans is 1 or -1.4

OpenStudy (anonymous):

wait, how do you know to put -2 +- sqrt 144 over 10? @meinme

OpenStudy (anonymous):

its formula

OpenStudy (anonymous):

okay, i just dont understand how you get those numbers, from the numbers i gave you? does that make sense,

OpenStudy (anonymous):

yeah do u know the formula of finding x in quadratic?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

{-b+-sqrt(b^2-4ac}/2a

OpenStudy (anonymous):

right, so how do you determine which is a, b, and c...

OpenStudy (anonymous):

general quadratic equation is ax^2+bx+c=0 now compare your equation youll get a, b and c

OpenStudy (anonymous):

Ohh, okaay, i see. Thank you so much.